# Better Estimation of the KL Divergence Between Language Models

### v2

@@ Abstract
Estimating the Kullback–Leibler (KL) divergence between language models has many applications, e.g., reinforcement learning from human feedback (RLHF), interpretability, and knowledge distillation. However, computing the exact KL divergence between two arbitrary language models is intractable. Thus, practitioners often resort to the use of sampling-based estimators. While it is easy to fashion a simple Monte Carlo (MC) estimator that provides an unbiased estimate of the KL divergence between language models, this estimator notoriously suffers from high variance, and can even result in a negative estimate of the KL divergence, a non-negative quantity. In this paper, we introduce a Rao–Blackwellized estimator that is also unbiased and provably has variance less than or equal to that of the standard Monte Carlo estimator. In an empirical study on sentiment-controlled fine-tuning, we show that our estimator provides more stable KL estimates and reduces variance substantially in practice. Additionally, we derive an analogous Rao–Blackwellized estimator of the gradient of the KL divergence, which leads to more stable training and produces models that more frequently appear on the Pareto frontier of reward vs. KL compared to the ones trained with the MC estimator of the gradient.
                     [Uncaptioned image] https://github.com/rycolab/kl-rb
@@ 1 Introduction
The Kullback–Leibler (KL; Kullback and Leibler, 1951) divergence is a statistical divergence that quantifies how one probability distribution differs from another. Measuring the KL divergence between probability distributions is a well-established problem that has been studied extensively in the statistics literature (Csiszár, 1967; Gibbs and Su, 2002, inter alia). In some special cases, e.g., in the case that we wish to measure the KL divergence between two Gaussian measures, the KL divergence has an analytical solution. However, in the general case, exact computation of the KL divergence is not analytically tractable or approximable with an efficient algorithm (Hershey and Olsen, 2007). This paper treats the case of computing the KL divergence between two language models (LMs), [^1] In the general case, computing the KL divergence between two arbitrary LMs is still hard. See Cortes et al. (2008) for a proof of the hardness of computing the KL between two arbitrary non-deterministic finite-state language models, a relatively well-behaved family of LMs. a fundamental task in natural language processing with numerous practical applications.
  • In reinforcement learning from human feedback (RLHF Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022) the KL divergence serves as a regularization term that constrains the distance between the language model and a reference model, typically the language model prior to fine-tuning. Specifically, constraining the KL divergence from the reference model prevents reward over-optimization and ensures that the model retains fluency while incorporating feedback from humans. Importantly, optimizing the RLHF objective needs more than just estimating the KL divergence; we also need to estimate its gradient with respect to the model’s parameters.
  • In research on the interpretability of neural language models, the KL divergence is used to quantify how a specific prompt affects the language model conditioned on the prompt, e.g., researchers have applied the KL divergence between the model’s distribution before and after prompting the model in a controlled fashion (Pezeshkpour, 2023; Du et al., 2024a; Stoehr et al., 2024).
  • As an evaluation metric, the KL divergence between the two language models serves as a metric for how well a neural language model has learned a target distribution. For example, Borenstein et al. (2024) and Svete et al. (2024) assess the learnability of probabilistic finite-state automata by measuring the KL divergence between a neural language model and the target distribution defined by the automata. Another example is evaluating language models after fine-tuning for human preference alignment, where we expect the model to stay close to the reference model (Rafailov et al., 2023).
  • In knowledge distillation, when training a student language model to approximate a teacher model, minimizing the divergence between their distribution is a common objective, e.g., Agarwal et al. (2024) optimize a token-level application of the KL divergence for this purpose.
The points above demonstrate that measuring the KL divergence between two language models is useful and widespread. However, in the case of neural language models, it is far from straightforward. It is easy to see why: Given an alphabet of symbols $\Sigma$ and two language models $p$ and $q$, distributions over $\Sigma^{*}$, the KL divergence is given by the following expression
@@
$\textsc{kl}(p\mid\mid q)=\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}\geq 0.$ (1)
@@
Recalling that $\Sigma^{*}$ is a countably infinite set, we cannot expect, in general, to compute Eq. 1 exactly in finite time without additional assumptions. While in some very special cases, e.g., where $p$ and $q$ are deterministic finite-state automata, there exist efficient algorithms, (Lehmann, 1977; Li and Eisner, 2009; Borenstein et al., 2024), we should not expect such an algorithm to exist in the case where $p$ and $q$ are neural language models, e.g., those based on the transformer (Vaswani et al., 2017; Radford et al., 2019; Meta, 2023). Thus, most researchers turn to approximation. A popular form of approximation resorts to stochastic approximation through Monte Carlo.
@@
The naive Monte Carlo estimator for KL divergence (Eq. 1) involves sampling $M$ strings $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$ and then averaging $\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}$. Even though this estimator is unbiased, it often exhibits high variance. More pathologically, the naive Monte Carlo estimator can result in negative estimates of KL, which may be undesirable in practice. To address these issues, practitioners adopt alternative techniques to ensure non-negativity. For example, Schulman (2020) proposes an unbiased, non-negative KL estimator that is widely used in practice (von Werra et al., 2020; Havrilla et al., 2023). However, it offers no theoretical variance improvement over the Monte Carlo estimator and, as we show empirically, can exhibit unbounded variance.
@@
In this paper, we derive an improved estimator of KL using Rao–Blackwellization (Gelfand and Smith, 1990; Casella and Robert, 1996), a well-established variance reduction technique from statistics. [^2] Despite its simplicity, our proposed estimator is absent from existing literature and open-source RLHF libraries von Werra et al. (2020); Havrilla et al. (2023); Hu et al. (2024); Sheng et al. (2024), highlighting a gap we believe is worth addressing. This results in an estimator that is provably unbiased and has a variance that is always less than or equal to that of the standard Monte Carlo estimator, while requiring no additional computational overhead. As a point of comparison to our Rao–Blackwellized estimator, we also provide a comprehensive formal analysis of various existing methods for estimating KL divergence, examining their bias and variance.
@@
We empirically validate our theoretical findings using the well-established sentiment-controlled generation task (Rafailov et al., 2023) as a testbed. Specifically, we measure the KL divergence between a GPT-2 model (Radford et al., 2019) before and after fine-tuning, where the fine-tuning objective is to steer the model toward generating positive movie reviews. Our experimental results confirm that the Rao–Blackwellized estimator significantly reduces the variance of the Monte Carlo estimator, yielding the most stable and reliable estimates among all methods studied. In contrast, alternative estimators from the literature fail to achieve meaningful variance reduction, and in some cases, lead to unbounded variance. We further examine how using our derived estimator in the fine-tuning loop of RLHF impacts the downstream performance. Our results suggest that using our Rao–Blackwellized estimator reduces the instability across different RLHF runs. We further look at the Pareto frontier of average rewards achieved by the model vs. its KL divergence with the reference model. We observe that models fine-tuned using the Rao–Blackwellized estimator appear significantly more often on the Pareto frontier of reward vs. KL compared to the models fine-tuned with the MC estimator.
@@ 2 Preliminaries
2.1 Language Modeling
Let $\Sigma$ be an alphabet, a finite, non-empty set of symbols. A string is a finite sequence of symbols drawn from the alphabet. The Kleene closure of an alphabet $\Sigma^{*}$ is the set of all strings with symbols drawn from $\Sigma$. A language model $p$ is a distribution over $\Sigma^{*}$. The prefix probability function $\vec{p}$ of a prefix $\boldsymbol{x}\in\Sigma^{*}$ is
@@ 2.1 Language Modeling
which is the cumulative probability of all strings in the language that have $\boldsymbol{x}$ as their prefix. We denote the conditional prefix probability as follows
$\vec{p}(\boldsymbol{x}\mid\boldsymbol{y})=\frac{\vec{p}(\boldsymbol{y}\boldsymbol{x})}{\vec{p}(\boldsymbol{y})}$ (3)
where we additionally define $\vec{p}(\textsc{eos}\mid\boldsymbol{y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\frac{p(\boldsymbol{y})}{\vec{p}(\boldsymbol{y})}$.
Language models can be factored into the product of distributions using the chain rule of probability, i.e., for any string $\boldsymbol{y}=y_{1}\cdots y_{N}\in\Sigma^{*}$ we can write
@@
where $\boldsymbol{y}_{<n}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}y_{1}\cdots y_{n-1}$ and $\textsc{eos}\notin\Sigma$ is a distinguished end-of-string symbol. Let $\overline{\Sigma}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\Sigma\cup\{\textsc{eos}\}$. In Eq. 4, each $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ can fruitfully be viewed as a distribution over $\overline{\Sigma}$. Despite the overloading of the notation, whether $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ refers to a prefix probability, a function which takes an argument from $\Sigma^{*}$, or a distribution over $\overline{\Sigma}$ will always be clear from context. Throughout this paper, we use $\boldsymbol{Y}$ to represent the string-valued random variable sampled from $p$. When taking $M$ samples from $p$, we use $\boldsymbol{Y}^{(m)}$ to denote the $m^{\text{th}}$ sample.
Let $p$ and $q$ be language models over a common alphabet $\Sigma$. The KL divergence between $p$ and $q$ is defined as follows: [^3] Whenever $p(\boldsymbol{y})$ is zero, the corresponding KL term is interpreted as zero because $\lim_{x\to 0^{+}}x\log x=0$.
$\textsc{kl}(p\mid\mid q)\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}=\operatorname*{\mathbb{E}}\left[f(\boldsymbol{Y})\right],$ (5)
where the function that we want to estimate is denoted with $f$, which in this case is $\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$. Indeed, for the remainder of the paper, we define $f(\boldsymbol{Y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$. Note that while the exact KL value is always non-negative, $f(\boldsymbol{Y})$ may be positive or negative. Importantly, the exact computation of Eq. 5 involves iterating over all strings in $\Sigma^{*}$.
In conditional generation tasks, such as dialogue generation or summarization, language models are typically prompted with an input string $\boldsymbol{x}\in\Sigma^{*}$. Therefore, the prompted language model defines a conditional distribution over strings given a prompt $\boldsymbol{x}$ denoted as $p(\cdot\mid\boldsymbol{x})$. In such cases, the average KL divergence between the language models over a corpus of prompts is reported. For notational simplicity, however, we omit the explicit conditioning on $\boldsymbol{x}$ and simply write $p(\boldsymbol{y})$. The results presented here generalize to the conditional setting.
2.2 Monte Carlo Estimation of KL
@@
The most common way to estimate the KL divergence is with the Monte Carlo estimator defined as
$\delta_{\textsc{mc}}$ $=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}=\frac{1}{M}\sum_{m=1}^{M}f(\boldsymbol{Y}^{(m)}),$ (6a)
@@
where $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$. Throughout the paper, we assume that the KL divergence is finite, i.e., $\textsc{kl}(p\mid\mid q)<\infty$. It is straightforward to show $\delta_{\textsc{mc}}$ is unbiased, i.e., $\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}]=\textsc{kl}(p\mid\mid q)$. The variance of this estimator is
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}}\right]$ $=\frac{1}{M^{2}}\mathop{\mathrm{Var}}\left[{\normalcolor\sum_{m=1}^{M}f(\boldsymbol{Y}^{(m)})}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f(\boldsymbol{Y})}\right].$ (7a)
@@
The true KL divergence is always non-negative, but Monte Carlo estimates can sometimes be negative. This happens because the estimate is based on a limited number of samples, and some sample draws can lead to negative values. This can be problematic during RLHF that depends on the KL divergence being non-negative.
In App. A we discuss the Horvitz–Thompson estimator, which is another unbiased estimator of the KL divergence.
2.3 Control Variate Monte Carlo

### v3

# Better Estimation of the Kullback–Leibler Divergence Between Language Models

@@ Abstract
Estimating the Kullback–Leibler (KL) divergence between language models has many applications, e.g., reinforcement learning from human feedback (RLHF), interpretability, and knowledge distillation. However, computing the exact KL divergence between two arbitrary language models is intractable. Thus, practitioners often resort to sampling-based estimators. While it is easy to fashion a simple Monte Carlo (MC) estimator that provides an unbiased estimate of the KL divergence between language models, this estimator notoriously suffers from high variance and can even result in a negative estimate of the KL divergence, a non-negative quantity. In this paper, we introduce a Rao–Blackwellized estimator that is unbiased and provably has variance less than or equal to that of the standard Monte Carlo estimator. In an empirical study on sentiment-controlled fine-tuning, we show that our estimator provides more stable KL estimates and reduces variance substantially. Additionally, we derive an analogous Rao–Blackwellized estimator of the gradient of the KL divergence, which leads to more stable training and produces models that more frequently appear on the Pareto frontier of reward vs. KL compared to the ones trained with the MC estimator of the gradient.
@@ 1 Introduction
The Kullback–Leibler [KL; 19] divergence is a statistical divergence that quantifies how one probability distribution differs from another. Measuring the KL divergence between probability distributions is a well-established problem that has been studied extensively in the statistics literature [7, 12, inter alia]. In some special cases, e.g., in the case that we wish to measure the KL divergence between two Gaussian measures, the KL divergence has an analytical solution. However, in the general case, exact computation of the KL divergence is not analytically tractable or approximable with an efficient algorithm [14]. This paper treats the case of computing the KL divergence between two language models (LMs), a fundamental task in natural language processing with numerous practical applications.
The KL divergence plays a central role across multiple applications. In reinforcement learning from human feedback [RLHF; 6, 35, 26], it is used as a regularization term to constrain the fine-tuned model from drifting too far from a reference model, preserving fluency and preventing reward over-optimization. In interpretability research, KL divergence quantifies how a specific prompt shifts the model distribution by comparing the model’s distributions before and after controlled interventions [27, 9, 36]. As an evaluation metric, KL divergence is used to assess how well language models approximate target distributions [4, 37]. In knowledge distillation, KL divergence is minimized to align a student model with a teacher model [1].
The above applications demonstrate that measuring the KL divergence between two language models is useful and widespread. However, in the case of neural language models, it is far from straightforward. It is easy to see why: Given an alphabet of symbols $\Sigma$ and two language models $p$ and $q$, distributions over $\Sigma^{*}$, the KL divergence is given by the following expression: [^1] Throughout this paper $\log$ denotes the natural logarithm function; thus, KL divergence is measure in nats rather than bits. We also note that terms of the form $p(\boldsymbol{y})\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}$ in Eq. 1 where $p(\boldsymbol{y})=0$ can correctly be taken to equal zero because $\lim_{p\to 0^{+}}p\log p/q=0$., [^2] In conditional tasks like dialogue generation, language models are prompted with an input $\boldsymbol{x}\in\Sigma^{*}$, inducing a conditional distribution $p(\cdot\mid\boldsymbol{x})$. KL divergences are typically averaged over a set of prompts. For simplicity, we omit $\boldsymbol{x}$ in notation and write $p(\boldsymbol{y})$; all of our results extend straightforwardly to the conditional case.
@@
$\textsc{kl}(p\mid\mid q)\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}.$ (1)
@@
Recalling that $\Sigma^{*}$ is a countably infinite set, we cannot expect, in general, to compute Eq. 1 exactly in finite time without additional assumptions. [^3] In general, computing the KL divergence between two arbitrary LMs exactly is undecidable. To see why, assume that each of the two language models is a probabilistic context-free grammar. In this case, deciding whether their KL divergence is zero is undecidable, as it follows directly from the undecidability of testing equivalence between two unweighted context-free grammars [15]. In the more restrictive case of probabilistic finite-state language models, it is pspace-hard. Importantly, however, the intractablity of exact computation does not imply that approximation is intractable. In practice, one can often obtain good Monte Carlo estimates of the KL divergence, provided that its variance is well-controlled. We study very practical methods for improving variance in this paper. While in some very special cases, e.g., where $p$ and $q$ are deterministic finite-state automata, there exist efficient algorithms, [20, 22, 4], we should not expect such an algorithm to exist in the case where $p$ and $q$ are neural language models, e.g., those based on the transformer [39, 28, 25]. Thus, most researchers turn to approximation, with Monte Carlo estimation being the most widely used method.
@@
The Monte Carlo (MC) estimator for KL divergence (Eq. 1) involves sampling $M$ strings $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$ and then averaging $\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}$. Even though this estimator is unbiased, it often exhibits high variance, which means the approximation can be noisy and unreliable. [^4] Monte Carlo estimation formally requires that the underlying random variable have finite variance; if the variance is unbounded, the estimator no longer converges in the limit. More pathologically, the naive MC estimator can result in negative estimates of KL, which may be undesirable in practice. [^5] Since the KL divergence is non-negative by definition, a negative estimate can be problematic when KL is used as part of a loss function, as it may destabilize the learning dynamics. To address these issues, practitioners adopt alternative techniques to ensure non-negativity. For example, Schulman [33] proposes an unbiased, non-negative KL estimator that is widely used in practice [40, 13]. However, the proposed method, in its original form, does not theoretically yield an estimator with lower variance, and, as we show empirically, can exhibit enormous variance.
@@
In this paper, we derive an improved estimator of KL using Rao–Blackwellization [RB; 30, 3, 11, 5], a well-established variance-reduction technique from statistics. [^6] Despite its simplicity, our proposed estimator is absent from existing literature and open-source RLHF libraries [40, 13, 17, 34], highlighting a gap we believe is worth addressing. This results in an estimator that is provably unbiased and has a variance that is always less than or equal to that of the standard Monte Carlo estimator, while requiring no additional computational overhead. As a point of comparison to our RB estimator, we also provide a comprehensive formal analysis of various existing methods for estimating KL divergence, examining their bias and variance.
@@
We empirically validate our theoretical findings using the sentiment-controlled generation task [29] as a testbed. Specifically, we measure the KL divergence between a GPT-2 model [28] before and after fine-tuning, where the fine-tuning objective is to steer the model toward generating positive movie reviews. Our experimental results confirm that our proposed estimator significantly reduces the variance of the Monte Carlo estimator, yielding the most stable and reliable estimates among all methods studied. In contrast, alternative estimators from the literature fail to achieve meaningful variance reduction, and in some cases, lead to unbounded variance. We further examine how using our derived estimator in the fine-tuning loop of RLHF impacts the downstream performance. Our results suggest that using our Rao–Blackwellized estimator reduces the instability across different RLHF runs. We further look at the Pareto frontier of average rewards achieved by the model vs. its KL divergence with the reference model. We observe that models fine-tuned using the RB estimator appear significantly more often on the Pareto frontier of reward vs. KL compared to the models fine-tuned with the MC estimator.
@@ 2 Preliminaries
2.1 Language Models
Let $\Sigma$ be an alphabet, a finite, non-empty set of symbols. A string is a finite sequence of symbols from $\Sigma$. Let $\Sigma^{*}$ denote the set of all such strings. A language model $p$ is a distribution over $\Sigma^{*}$. The prefix probability function $\vec{p}$ of a prefix $\boldsymbol{x}\in\Sigma^{*}$ is
@@ 2.1 Language Models
which is the cumulative probability of all strings in the language that have $\boldsymbol{x}$ as their prefix. We denote the conditional prefix probability as $\vec{p}(\boldsymbol{y}\mid\boldsymbol{x})=\frac{\vec{p}(\boldsymbol{x}\boldsymbol{y})}{\vec{p}(\boldsymbol{x})}$, where we additionally define $\vec{p}(\textsc{eos}\mid\boldsymbol{y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\frac{p(\boldsymbol{y})}{\vec{p}(\boldsymbol{y})}$. Language models can be factored into the product of distributions using the chain rule of probability, i.e., for any string $\boldsymbol{y}=y_{1}\cdots y_{N}\in\Sigma^{*}$ we can write
@@
where $\boldsymbol{y}_{<n}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}y_{1}\cdots y_{n-1}$ and $\textsc{eos}\notin\Sigma$ is a distinguished end-of-string symbol. Let $\overline{\Sigma}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\Sigma\cup\{\textsc{eos}\}$. In Eq. 3, each $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ can fruitfully be viewed as a distribution over $\overline{\Sigma}$. Despite the overloading of the notation, whether $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ refers to a prefix probability, a function which takes an argument from $\Sigma^{*}$, or a distribution over $\overline{\Sigma}$ will always be clear from context. Throughout this paper, we use $\boldsymbol{Y}$ to represent the string-valued random variable sampled from pp. When taking $M$ i.i.d. samples from $p$, we use $\boldsymbol{Y}^{(m)}$ to denote the $m^{\text{th}}$ sample.
2.2 Monte Carlo KL Estimation
@@
A simple way to estimate the KL divergence is with the Monte Carlo estimator (MC) defined as
$\mu_{\textsc{mc}}$ $=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}=\frac{1}{M}\sum_{m=1}^{M}f(\boldsymbol{Y}^{(m)}),$ (4)
@@
where $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$ and $f(\boldsymbol{Y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$. Throughout the paper, we assume that the KL divergence is finite, i.e., $\textsc{kl}(p\mid\mid q)<\infty$. It is straightforward to show $\mu_{\textsc{mc}}$ is unbiased, i.e., $\operatorname*{\mathbb{E}}[\mu_{\textsc{mc}}]=\textsc{kl}(p\mid\mid q)$ and the variance of this estimator is $\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{mc}}}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f(\boldsymbol{Y})}\right]$. In App. A, we discuss the Horvitz–Thompson estimator, another unbiased KL estimator.
@@
Note that while the exact KL value is always non-negative, $f(\boldsymbol{Y})$ may be positive or negative. Consequently, the Monte Carlo estimate $\mu_{\textsc{mc}}$ may also be negative. This happens because the estimate is based on a limited number of samples, and some sample draws can lead to negative values. This can be problematic during RLHF, which depends on the KL divergence being non-negative.
2.3 Control Variate Monte Carlo Estimation